If each of the points (x1,4) , (-2,y1) lies on the line joining the points (2,-1),(5,-3) then the point P(x1,y1) lies on the lines ?
Give the equation of that line.

Grade:12

2 Answers

Nishant Vora IIT Patna
9 years ago
equation of line joining (2,-1),(5,-3) is:
(y+1)=\frac{-1+3}{2-5}(x-2)
(y+1)=\frac{2}{-3}(x-2)
2x + 3y -1 = 0

Now (x1,4) , (-2,y1) lie on 2x + 3y -1 = 0. So it will satisfy.
2x1 + 3*4 -1 = 0 and 2*-2 + 3y1 -1 = 0
x1=-11/2 and y1 = 5/3

Now 2(-11/2) + 3(5/3) -1 does not= 0
So (x1,y1) does not lie on this line.


Aditya Pati
9 years ago
Sir,Actually they are asking on which line x1,y1 lies :) ?